OpenStudy (anonymous):

Can someone help me find the slope of the line 5x - 2y = -10?

OpenStudy (australopithecus):

solve for y then use the equation y = mx + b where m is the slope and b is the y intercept

OpenStudy (australopithecus):

After you solve for y compare to the equation of the line to find the slope

OpenStudy (anonymous):

equation of a line is y=mx+b.. now to in order for our equation to agree with line equation, we have to rearrange it. 5x - 2y = -10 transfer -2y to right side, makes it +2y and -10 to left side, makes it +10 5x + 10 = 2y or 2y = 5x +10 we only need 1 y. so divide both sides by 2.. y = (5/2)x + 5 y = mx+b where m is the slope, b is the y-intercept.